# The Bernstein problem for Sobolev intrinsic graphs in the Heisenberg group

### v1

@@ 4. Consequence of the first variation
Let $f\in W^{1,q}_{\mathrm{loc}}(\omega)\cap C^{0}(\omega)$ with $q>2$ and such that (1) holds. Suppose that $f$ is stationary. Then there is $\hat{p}>2$ such that for every $p>\hat{p}$ the following holds: for every $(\eta_{0},\tau_{0})\in\omega$ there are $\epsilon>0$ and $\chi:(\eta_{0}-\epsilon,\eta_{0}+\epsilon)\times(\tau_{0}-\epsilon,\tau_{0}+\epsilon)\to\mathbb{R}$ such that the map $\Psi(t,\zeta)=(t,\chi(t,\zeta))$ is a $p$-bi-Sobolev $f$-Lagrangian homeomorphism and $\chi$ is of the form
@@
As a consequence, we have that $\nabla^{f}f\in C^{0}(\omega)$. The exponent $\hat{p}$ can be taken as in Lemma A.1.
@@
Take $\hat{p}$ from Lemma A.1 and $p>\hat{p}$. Take $\epsilon>0$ such that $\kappa=\frac{2\epsilon p^{2}}{p-1}$: then (1) and Theorem 3.9 imply the existence of such a $p$-bi-Sobolev $f$-Lagrangian homeomorphism $\Psi$. The form of $\chi$ and the regularity of $a,b$ come from Remark 4.2. The fact that $c(\zeta)=\zeta$ is a consequence of Theorem 3.9. The formulas $a(\zeta)=\nabla^{f}f(\eta_{0},\zeta)$ and $b(\zeta)=f(\eta_{0},\zeta)$ follow from (12). Finally, the continuity of $\nabla^{f}f$ follows from (12) and the fact that $(\eta_{0},\tau_{0})$ is arbitrary in $\omega$. ∎
@@ 5. Consequences of the second variation
Notice that, by Morrey Embedding Theorem, $f\in C^{0}(\mathbb{R}^{2})$. Being $f$ is stationary we can apply Corollary 4.3 at every point $(\eta_{0},\tau_{0})\in\mathbb{R}^{2}$. So, in a neighborhood of $(\eta_{0},\tau_{0})$, we obtain a $p$-bi-Sobolev $f$-Lagrangian homeomorphism of the form $\Psi(t,\zeta)=(t,\chi(t,\zeta))$ with
for some $a,b\in W^{1,p}_{\mathrm{loc}}((\tau_{0}-\epsilon,\tau_{0}+\epsilon))\cap C^{0}((\tau_{0}-\epsilon,\tau_{0}+\epsilon))$.
In particular, from Definition 3.7, we have the formula
$f\left(t,a(\zeta)\frac{t^{2}}{2}+b(\zeta)t+\zeta\right)=a(\zeta)(t-\eta_{0})+b(\zeta).$
By Lemma A.1, the hypothesis $q>4$ ensures that we can choose $p$ so that both (23) and (25) are verified, and thus we can apply Proposition 5.7. We obtain that $a^{\prime}=b^{\prime}=0$.
In conclusion, we have that, every $(\eta_{0},\tau_{0})\in\mathbb{R}^{2}$ has a neighborhood where $f(\eta,\tau)=a(\eta-\eta_{0})+b$ for some $a,b\in\mathbb{R}$ depending on $(\eta_{0},\tau_{0})$. The only functions on $\mathbb{R}^{2}$ with this property are affine functions. Indeed, this implies that $f\in C^{\infty}(\mathbb{R}^{2})$ and $\frac{\partial f}{\partial\tau}=\frac{\partial^{2}f}{\partial\eta^{2}}=0$ in $\mathbb{R}^{2}$. Therefore, it follows that there exist $a,b\in\mathbb{R}$ with $f(\eta,\tau)=a\eta+b$ for all $\eta,\tau\in\mathbb{R}$. ∎

### v2

@@ 4. Consequence of the first variation
Let $f\in W^{1,q}_{\mathrm{loc}}(\omega)\cap C^{0}(\omega)$ with $q>2$ and such that (1) holds. Suppose that $f$ is stationary. Then for every $p\geq 1$ the following holds: for every $(\eta_{0},\tau_{0})\in\omega$ there are $\epsilon>0$ and $\chi:(\eta_{0}-\epsilon,\eta_{0}+\epsilon)\times(\tau_{0}-\epsilon,\tau_{0}+\epsilon)\to\mathbb{R}$ such that the map $\Psi(t,\zeta)=(t,\chi(t,\zeta))$ is a $p$-bi-Sobolev $f$-Lagrangian homeomorphism, where $\chi$ is of the form
@@
As a consequence, we have that $f,\nabla^{f}f\in C^{0}(\omega)\cap W^{1,p}_{\mathrm{loc}}(\omega)$, for every $p\geq 1$. Moreover, if $\omega=\mathbb{R}^{2}$, then we can take $\epsilon=\infty$; in particular, the function $\Psi(t,\zeta)=(t,\chi(t,\zeta))$ with
(23) $\chi(t,\zeta)=\frac{\nabla^{f}f(0,\zeta)}{2}t^{2}+f(0,\zeta)t+\zeta$
is a $p$-bi-Sobolev $f$-Lagrangian homeomorphism $\mathbb{R}^{2}\to\mathbb{R}^{2}$, for every $p\geq 1$.
@@
Let $p\geq 1$. Since the condition of being $p$-bi-Sobolev is local, we assume $p>\hat{p}$, where $\hat{p}$ is chosen as in Lemma A.1. Take $\epsilon>0$ such that $\kappa=\frac{2\epsilon p^{2}}{p-1}$: then (1) and Theorem 3.9 imply the existence of such a $p$-bi-Sobolev $f$-Lagrangian homeomorphism $\Psi$. The form of $\chi$ and the regularity of $a,b$ come from Remark 4.2. The fact that $c(\zeta)=\zeta$ is a consequence of Theorem 3.9. The formulas $a(\zeta)=\nabla^{f}f(\eta_{0},\zeta)$ and $b(\zeta)=f(\eta_{0},\zeta)$ follow from (12). This shows the first part of the corollary.
The continuity of $\nabla^{f}f$ follows from (12) and the fact that $(\eta_{0},\tau_{0})$ is arbitrary in $\omega$. The Sobolev regularity of $f$ and $\nabla^{f}f$ is more tricky to show, but not difficult: First, notice that, in a neighborhood of $(\eta_{0},\tau_{0})$, we have $f=\partial_{t}\chi\circ\Psi^{-1}=(at+b)\circ\Psi^{-1}$ and $\nabla^{f}f=\partial_{t}^{2}\chi\circ\Psi^{-1}=a\circ\Psi^{-1}$, where $a$ and $b$ belong to the class $W^{1,p}_{\mathrm{loc}}$. Second, using the first part of the corollary, we notice that, up to taking a smaller neighborhood of $(\eta_{0},\tau_{0})$, for $\bar{q}$ arbitrarily large, we have that $\Psi$ is $\bar{q}$-bi-Sobolev homeomorphism and both $a$ and $b$ are in the class $W^{1,\bar{q}}_{\mathrm{loc}}$. Third, we apply Proposition 3.4 with $\bar{p}$ large enough, $\bar{s}:=\bar{p}/2$ and $\bar{q}$ given by the formula $\bar{q}:=\frac{\bar{p}^{2}\bar{s}}{(\bar{p}-\bar{s})(\bar{p}-2)}$: we only need $\bar{p}$ to be large enough to have $\bar{q}\geq 1$. Note that, since $\bar{q}>\bar{p}$, $\Psi$ is a $\bar{p}$-bi-Sobolev homeomorphism too. So, from Proposition 3.4 we obtain that $f$ and $\nabla^{f}f$ belong to $W^{1,\bar{s}}_{\mathrm{loc}}(U)$, for some neighborhood $U$ of $(\eta_{0},\tau_{0})$. Since we can take $\bar{s}$ larger than $p$, so that we conclude $f,\nabla^{f}f\in W^{1,\bar{s}}_{\mathrm{loc}}(U)$. Since $(\eta_{0},\tau_{0})$ is arbitrary, we conclude $f,\nabla^{f}f\in W^{1,p}_{\mathrm{loc}}(\omega)$.
Next, suppose $\omega=\mathbb{R}^{2}$ and fix $(\eta_{0},\tau_{0})\in\mathbb{R}^{2}$. Without loss of generality, we assume $(\eta_{0},\tau_{0})=(0,0)$, so that $\Psi(t,\zeta)=(t,\chi(t,\zeta))$ with $\chi$ as in (23). The fact that the map $\Psi$ is a homeomorphism $\mathbb{R}^{2}\to\mathbb{R}^{2}$ is proven in [12, Lemmas 3.3 and 3.5]. We need to show that $\Psi,\Psi^{-1}\in W^{1,p}_{\mathrm{loc}}(\mathbb{R}^{2};\mathbb{R}^{2})$, each $p\geq 1$. Since $a,b\in W^{1,p}_{\mathrm{loc}}(\mathbb{R})$, then $\Psi\in W^{1,p}_{\mathrm{loc}}(\mathbb{R}^{2};\mathbb{R}^{2})$. For the inverse map, notice that, if $(t,\zeta),(\eta,\tau)\in\mathbb{R}^{2}$ are such that $(\eta,\tau)=\Psi(t,\zeta)$, then $\eta=t$, $\tau=\chi(t,\zeta)=a(\zeta)t^{2}/2+b(\zeta)t+\zeta$, $f(\eta,\tau)=f(t,\chi(t,\zeta))=\partial_{t}\chi(t,\zeta)=a(\zeta)t+b(\zeta)$, and $\nabla^{f}f(\eta,\tau)=\nabla^{f}f(t,\chi(t,\zeta))=\partial_{t}^{2}\chi(t,\zeta)=a(\zeta)$. Therefore,
$\tau$ $=a(\zeta)t^{2}/2+b(\zeta)t+\zeta$
$=\nabla^{f}f(\eta,\tau)t^{2}/2+(f(\eta,\tau)-a(\zeta)t)t+\zeta$
$=-\nabla^{f}f(\eta,\tau)t^{2}/2+f(\eta,\tau)t+\zeta,$
that is,
$\zeta=\tau+\nabla^{f}f(\eta,\tau)\eta^{2}/2-f(\eta,\tau)\eta.$
We conclude that $\Psi^{-1}(\eta,\tau)=(\eta,\tau+\nabla^{f}f(\eta,\tau)\eta^{2}/2-f(\eta,\tau)\eta)$. Since $f,\nabla^{f}f\in W^{1,p}_{\mathrm{loc}}(\mathbb{R}^{2})$, we obtain $\Psi^{-1}\in W^{1,p}_{\mathrm{loc}}(\mathbb{R}^{2};\mathbb{R}^{2})$. ∎
@@ 5. Consequences of the second variation
Notice that, by Morrey Embedding Theorem, $f\in C^{0}(\mathbb{R}^{2})$. Being $f$ is stationary, we apply Corollary 4.3: we obtain a $f$-Lagrangian homeomorphism $\Psi:\mathbb{R}^{2}\to\mathbb{R}^{2}$ of the form $\Psi(t,\zeta)=(t,\chi(t,\zeta))$ with
$\chi(t,\zeta)=\frac{a(\zeta)}{2}t^{2}+b(\zeta)t+\zeta,$
for some $a,b\in W^{1,p}_{\mathrm{loc}}(\mathbb{R})\cap C^{0}(\mathbb{R})$. Moreover, the map $\Psi$ is $p$-bi-Sobolev for every $p\geq 1$. By Lemma A.1, the hypothesis $q>4$ ensures that we can choose $p$ so that both (24) and (26) are verified, and thus we can apply Proposition 5.7. We obtain that $a^{\prime}=b^{\prime}=0$. We conclude that $f(\eta,\tau)=a\eta+b$ is affine. ∎